Hello, I am working on the problem 9x + 3y - 5 = 0 and 3x + y = 5/3. It's asking to solve the system of equations using substitution. I don't get what I have to plug in to get the answer?
Answer by mananth(16946)

9 x + 3 y = 5 .............1
1 y = 1.67 -3 x ---------2
/ 1 =
y = 1.67 -3 x
Substitute y in (1)
9x+3(1.67-3x)= 5
9x+5+-9 x=5
9x-9x=5 -5
0x=0
There is no solution to this system
they are parallel lines with different y intercepts slopes are same